Problem 3

All equity company X trades at a P/E multiple of 25 and is priced correctly under Gordon’s model. It has a beta of 1.2. Risk-free rate of return is 3% and the return on the market portfolio is 14%. The company pays out 40% of its earnings as dividends. The expected dividend next year is $1.50 per share. Find ROE.
Answer #1

D1 = $1.50

EPS1 = 1.50/0.40

EPS1 = $3.75

Price = 25(3.75) = $93.75

Using CAPM Model,

Cost of equity = 0.03 + 1.20(0.14 - 0.03)

Cost of equity = 16.20%

Using Constant Growth Model,

g = 0.1620 - 1.50/93.75

g = 14.60%

0.1460 = (1 - 0.40)ROE

ROE = 24.33%